OpenStudy (hellokitty17):

Functions f(x) and g(x) are shown below: f(x) = x2 g(x) = x2 + 8x + 16 In which direction and by how many units should f(x) be shifted to obtain g(x)? Left by 4 units Right by 4 units Left by 8 units Right by 8 units

OpenStudy (hellokitty17):

Explain the answer plz

oregonduck (oregonduck):

b

OpenStudy (hellokitty17):

Explain why please

OpenStudy (anonymous):

I think it shifts the right by 4 units. (a-b)^2=a^2-2ab+b^2 x^2 - 8x + 16 (same with a=x and b=4) x^2 - 8x + 16= (x-4)^2 So the shift from x^2 is by 4 units.